OpenStudy (anonymous):

Bill is able to save $35/week after working part-time and paying his expenses. These two formulas show his weekly savings: A. f(1) = 35, f(n) = f(1) + f(n-1) for n>1 B. f(n) = 35n II. Use the recursive formula to make a table of values for 1_< n _<5. show your calculations. Explain what your table means.

OpenStudy (amoodarya):

f(1) = 35, f(n) = f(1) + f(n-1) f(2)=f(1) +f(2-1)=2f(1)=2*35 f(3)=f(1)+f(3-1)=f(1)+f(2)=35+2*35 =3*35 .. so f(n) =f(1) +f(f-1)=35+(n-1)*35 =n*35